OpenStudy (anonymous):

Relationship B has a lesser rate than Relationship A. The graph represents Relationship A. Which table could represent Relationship B? First-quadrant graph showing a ray from the origin through the points (5, 2) and (10, 4). Horizontal axis label is Time in weeks. Vertical axis label is Plant growth in inches. A. Time (weeks) 3 4 6 9 Plant growth (in.) 1.8 2.4 3.6 5.4 B. Time (weeks) 3 4 6 9 Plant growth (in.) 0.9 1.2 1.8 2.7 C. Time (weeks) 3 4 6 9 Plant growth (in.) 1.5 2 3 4.5 D. Time (weeks) 3 4 6 9 Plant growth (in.) 2.7 3.6 5.4 8.1

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Angel_Halo23

OpenStudy (anonymous):

uhg math, i can try but no promises

OpenStudy (anonymous):

Ok i same all the time

OpenStudy (igreen):

Where's the graph for Relationship A?

OpenStudy (igreen):

Is this also multiple choice?

OpenStudy (anonymous):

I think c and no its not

OpenStudy (igreen):

Oh, then A is incorrect..B seems correct..let me check C and D.

OpenStudy (anonymous):

id think c is right

OpenStudy (igreen):

Yeah, C is incorrect.

OpenStudy (anonymous):

im thinking d??

OpenStudy (anonymous):

yea cause thats the closest 2 more after this

OpenStudy (igreen):

D is also incorrect, it has the highest rate than the rest..the answer is B.

OpenStudy (anonymous):

ohhk

OpenStudy (texaschic101):

the rate is the slope... Relationship A slope is 2/5 which is 0.4 so Relationship B has to have a slope of less then 2/5 (less then 0.4) Lets check what igreen said...lets check B (3,0.9) and (4,1.2) slope = (y2 - y1) / (x2 - x1) slope = (1.2 - 0.9) / (4 - 3) slope = 0.3/1 or 0.3 yep...it is B

OpenStudy (anonymous):

Hi how do i add fans

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

will give medal